Solve the following 2 x 2 systems using substitutions
2x + 7y =9
9x + 4y =13

Answer by checkley75(3666)

2X+7Y=9, 2X=-7Y+9, X=-7Y/2+9/2, X=-3.5Y+4.5 NOW SUBSTITUTE IN THE OTHER EQUATION
9(-3.5Y+4.5)+4Y=13
-31.5Y+40.5+4Y=13
-27.5Y=13-40.5
-27.5Y=-27.5
Y=-27.5/-27.5
Y=1 ANSWER. NOW SOLVE FOR X
2X+7*1=9
2X=9-7
2X=2
X=2/2
X=1 ANSWER.
-------------------------------------------------------------------
THIS IS THE METHOD. NOW GIVE IT A SHOT ON YOUR OWN FOR THE REST.